# Transforming H&E images into IHC: A Variance-Penalized GAN for Precision Oncology

### v1

@@ VI Discussion
The challenge of accurately translating hematoxylin and eosin (H&E) stained tissue into immunohistochemistry (IHC) representations, particularly for HER2-positive (IHC 3+) cases, have impeded broader clinical deployment of generative models in digital pathology. Existing frameworks often struggle to preserve fine-grained morphological and phenotypic features critical for accurate biomarker assessment. Our work addresses this long-standing limitation by introducing a novel variance-based penalty term into the pyramid pix2pix architecture, which fundamentally reshapes the loss penalty landscape to encourage diversity and fidelity in the generated images. The proposed modification to the loss function, detailed in eq. 10, effectively addresses the mode collapse problem by introducing a penalty for generating less diverse images. The results affirm the effectiveness of this modification across three pivotal dimensions: improved HER2-positive translation accuracy, mitigation of mode collapse, and generalizability to non-medical image translation tasks.
VI-A High-Fidelity Translation of HER2-Positive (IHC 3+) Images
@@
HER2-positive tumors, marked by strong circumferential membrane staining, present high morphological heterogeneity that has historically confounded generative models. While prior models such as pix2pix and pyramid pix2pix achieved reasonable performance on low-expression IHC classes (0/1+/2+), they consistently failed to generate realistic IHC 3+ images, a shortcoming with direct clinical implications. By incorporating a variance-based penalty that directly targets the model’s tendency toward oversimplified outputs, our method achieves substantial gains in PSNR and SSIM for IHC 3+ cases. Specifically, PSNR improved by over 9% and SSIM by nearly 20% relative to pyramid pix2pix, reflecting enhanced pixel-level fidelity and structural coherence. Crucially, the decrease in Fréchet Inception Distance (FID) for IHC 3+, a metric sensitive to perceptual realism and distributional alignment, demonstrates that our model not only improves quantitative performance but also generates images with higher diagnostic plausibility. As shown in Fig. 14, our method consistently produces more diverse and representative images across different classes of input, particularly in comparison to baseline methods such as pix2pix and pyramid pix2pix.
VI-B Variance-Based Penalty and Mitigation of Mode Collapse
@@
Mode collapse remains one of the most persistent and debilitating issues in GAN-based translation models, often manifesting as repetitive or degenerate outputs. Traditional solutions have explored architectural adjustments or adversarial training stabilization techniques, but few have directly targeted the diversity of outputs through the loss function. Our contribution departs from this paradigm by introducing a batch-level variance regularization term that penalizes the model when the variance in generated images deviates significantly from that of the real distribution. This variance-based constraint acts as a statistical anchor, encouraging the generator to produce outputs that span the intrinsic variability of the target domain.
@@
VI-C Generalizability Beyond Medical Imaging
The proposed model was initially optimized for histopathological image translation task, we further evaluated its generalization capability on the LLVIP dataset, a benchmark in the domain of visible-to-infrared translation. The ability of our model to outperform pyramid pix2pix in both PSNR and SSIM on this out-of-domain task highlights the robustness and adaptability of our variance-regularized framework. This observation is nontrivial: most domain-specific models overfit to localized feature distributions and exhibit poor transferability. Our results indicate that the variance-based penalty does not merely fine-tune the model for HER2 classification, but rather instills a broader inductive bias toward generative diversity, a quality that enhances performance even in unrelated modalities.
@@ VII Conclusion
VIII acknowledgment
The publication of this article was funded by Qatar National Library.

### v2

@@ 7 Discussion
The challenge of accurately translating hematoxylin and eosin (H&E) stained tissue into immunohistochemistry (IHC) representations has impeded broader clinical deployment of generative models in digital pathology. Existing frameworks often struggle to preserve fine-grained morphological and phenotypic features critical for accurate biomarker assessment. Our work addresses this long-standing limitation by introducing a novel variance-based penalty term into the pyramid pix2pix architecture, which fundamentally reshapes the loss penalty landscape to encourage diversity and fidelity in the generated images. The proposed modification to the loss function, detailed in eq. 11, effectively addresses the mode collapse problem by introducing a penalty for generating less diverse images. The results affirm the effectiveness of this modification across three pivotal dimensions: improved HER2-positive translation accuracy, mitigation of mode collapse, and generalizability to non-medical image translation tasks.
7.1 High-Fidelity Translation of HER2-Positive (IHC 3+) Images
@@
HER2-positive tumors, marked by strong circumferential membrane staining, present high morphological heterogeneity that has historically confounded generative models. While prior models such as pix2pix and pyramid pix2pix achieved reasonable performance on low-expression IHC classes (0/1+/2+), they consistently failed to generate realistic IHC 3+ images, a shortcoming with direct clinical implications. By incorporating a variance-based penalty that directly targets the model’s tendency toward oversimplified outputs, our method achieves substantial gains in PSNR and SSIM for IHC 3+ cases. Crucially, the 33% decrease in Fréchet Inception Distance (FID) for IHC 3+ against the state of the art methods, a metric sensitive to perceptual realism and distributional alignment, demonstrates that our model not only improves quantitative performance but also generates images with higher diagnostic plausibility. As shown in Figure 12, our method consistently produces more diverse and representative images across different classes of input, particularly in comparison to baseline methods such as pix2pix and pyramid pix2pix.
7.2 Variance-Based Penalty and Mitigation of Mode Collapse
@@
Mode collapse remains one of the most persistent and debilitating issues in GAN-based translation models, often manifesting itself as repetitive or degenerate outputs. Traditional solutions have explored architectural adjustments or adversarial training stabilization techniques, but few have directly targeted the diversity of outputs through the loss function. Our contribution departs from this paradigm by introducing a batch-level variance regularization term that penalizes the model when the variance in generated images deviates significantly from that of the real distribution. This variance-based constraint acts as a statistical anchor, encouraging the generator to produce outputs that span the intrinsic variability of the target domain.
Mode collapse arises when the generator fails to capture certain modes of the real data distribution. Two of the metrics that can quantify the reduction in mode collapse are Earth Mover’s Distance (EMD) (rubner2000earth) and Kullback–Leibler (KL) Divergence (kullback1951information). EMD quantifies the minimal effort needed to transform one distribution into another. The lower EMD for our model as shown in Figure 16(a) indicates that the distribution of generated images closely aligns with that of the ground truth. Similarly, Kullback–Leibler (KL) Divergence measures how one probability distribution diverges from a reference distribution. The lower KL divergence observed in Figure 16(b) for our model suggests a better approximation of the real data distribution and a reduction in mode collapse.
@@
Figure 16: Comparison of EMD and KL divergence across models. Boxplot (a) illustrates the distribution of EMD values, and boxplot (b) illustrates the distribution of KL divergence values computed between the color histograms of ground truth (GT) images and those generated by different models: Ours and Pyramid Pix2Pix. The lower median and tighter spread in the GT vs Ours distribution in both boxplots indicate higher similarity in pixel intensities and greater consistency across samples.
7.3 Generalizability Beyond Medical Imaging
The proposed model was initially optimized for histopathological image translation task, we further evaluated its generalization capability on the facades dataset (isola2017image), a benchmark in the domain of visible-to-infrared translation. The ability of our model to outperform pyramid pix2pix in PSNR, SSIM and FID on this out-of-domain task highlights the robustness and adaptability of our variance-regularized framework. This observation is nontrivial: most domain-specific models overfit to localized feature distributions and exhibit poor transferability. Our results indicate that the variance-based penalty does not merely fine-tune the model for HER2 classification, but rather instills a broader inductive bias toward generative diversity, a quality that enhances performance even in unrelated modalities.
@@ 8 Conclusion
9 Tables
Table 1: Comparison of Evaluation Metrics for pix2pix, pyramid pix2pix, and Our Method for IHC 0.
Metrics pix2pix pyramid pix2pix ours
PSNR 19.6 19.2 20.48
SSIM 0.52 0.46 0.55
FID 914.32 920.98 890.46
Table 2: Comparison of Evaluation Metrics for pix2pix, pyramid pix2pix, and Our Method for IHC 1+.
Metrics pix2pix pyramid pix2pix ours
PSNR 22.22 21.07 21.65
SSIM 0.47 0.44 0.48
FID 605.58 642.79 472.19
Table 3: Comparison of Evaluation Metrics for pix2pix, pyramid pix2pix, and Our Method for IHC 2+.
Metrics pix2pix pyramid pix2pix ours
PSNR 20.9 20.4 21.3
SSIM 0.44 0.42 0.46
FID 473.84 527.4 395.4
Table 4: Comparison of Evaluation Metrics for pix2pix, pyramid pix2pix, and Our Method for IHC 3+.
Metrics pix2pix pyramid pix2pix ours
PSNR 19.2 19.4 19.5
SSIM 0.39 0.38 0.42
FID 733.01 758.74 508.38
Table 5: Comparison of Evaluation Metrics for pix2pix, pyramid pix2pix and Our Method using all classes of BCI datatset
Metric pix2pix pyramid pix2pix ours
PSNR 20.74 21.15 22.16
SSIM 0.44 0.43 0.47
FID 472.6 516.75 346.37
Table 6: Comparison of evaluation metrics for Pyramid Pix2Pix and the proposed method on the Facades dataset.
Metric Pyramid Pix2Pix Ours
PSNR 12.91 12.80
SSIM 0.24 0.25
FID 611.34 549.07
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
SR (Sara Rehmat): Conceptualization, methodology, software development, data curation, formal analysis, validation, visualization, and writing, original draft preparation.
HR (Hafeez-ur-Rehman): Conceptualization, supervision, methodology refinement, investigation, resources, writing—review and editing, and project administration.
BGK (Byeong-Gwon Kang): Conceptualization, supervision, methodology guidance, validation, writing and editing, and funding acquisition.
SA (Sarra Ayouni): Data curation, investigation, validation, and writing, review and editing.
YN (Yunyoung Nam): Supervision, resources, writing, review and editing, validation and funding acquisition.
Funding
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2023-00218176) and the Soonchunhyang University Research Fund. Authors like to thanks Princess Nourah bint Abdulrahman University Researchers Supporting Project number (PNURSP2026R896), Princess Nourah bint Abdulrahman University, Riyadh,Saudi Arabia.
Acknowledgments
The authors acknowledge the Department of Computer Science, National University of Computer and Emerging Sciences (FAST-NUCES), Peshawar Campus, Pakistan, for providing hardware support and the necessary laboratory resources. We also thank the BCI dataset providers for making the histopathology images publicly available, which enabled this research. Computational resources were supported by Google Colaboratory (Google Colab).
Data Availability Statement
The datasets used and/or analyzed during the current study are available from the corresponding author upon request.